# Engagement-Driven Content Generation with Large Language Models

### v2

@@ Appendix B Engagement-Sentiment Correlation on Synthetic Network
Figure 8: Engagement resulting from $\mathcal{M}_{\epsilon}$ varying the message sentiment $s_{t}$. From top-left to bottom-right: opinion positive/negative/uniform/neutral/uniform; modularity high/low/high/low/high; homophily: high/low/low/high/low; position: central/comm-smallest/comm-largest/echo-high/echo-low.
Figure 8 depicts some examples of the engagement resulting from the propagation protocol $\mathcal{M}_{\epsilon}$ over the synthetic network, by varying the content sentiment $s_{t}$. From top-left to bottom-right: the first plot refers to a positely distributed network with high modularity and homophily and the starting position is central; the second one refers to a negatively distributed graph with low modularity and homophily and with Comm-smallest as source node; the third represents the engagement produced over a network with uniform opinion distribution, high modularity, low homophily and Comm-largest as starting node; the fourth plot refers to a neutral network with low modularity, high homophily and assumes the source node to be in a Echo-high position; and finally, the last plot represents the engagement over a uniformly distributed graph with high modularity, low homophily and the source node in the Echo-low configuration.
Appendix C Simulated/Actual Engagement on Real Network
@@
We validate the reliability of our propagation model by comparing the engagement estimated over the Brexit real network, and the actual number of retweets resulting from the most engaging content in the dataset. Figure 9 shows the corresponding regression line computed via Random Sample Consensus (RANSAC) (Cantzler, [n. d.]). Despite our propagation protocol tends to overestimate the engagement produced by the given content, there is a positive correlation between the number of active users estimated by our propagation model ($|\mathcal{A}|$) and the number of actual retweets, thus validating our findings over the real network.

### v4

@@ Appendix A Engagement-Sentiment Correlation
Figure 13: Engagement resulting from $\mathcal{M}_{\epsilon}$ varying the message sentiment $s_{t}$. From top-left to bottom-right: network’s opinion positive/negative/uniform/neutral/uniform; modularity high/low/high/low/high; homophily: high/low/low/high/low; position: central/comm-smallest/comm-largest/echo-high/echo-low.
The first plot corresponds to a network with positively distributed opinions, high modularity, and strong homophily, where the source node is centrally positioned. In this case, engagement grows non-monotonically, reaching higher values for strong positive opinions (i.e., sentiment). The second plot represents a negatively distributed network with low modularity and weak homophily, with the source node placed in the smallest community (Comm-smallest). Here, engagement decreases as opinions become more positive, following an inverse trend compared to the first case. The third plot depicts engagement in a network with a uniformly distributed opinion landscape, high modularity, and low homophily, where the source node is located in the largest community (Comm-largest). In this scenario, engagement remains at half of the maximum across a wide opinion range (0.15 to 0.85) and drops to zero elsewhere. The fourth plot corresponds to a neutral network with low modularity and high homophily, where the source node is positioned within a highly opinionated echo chamber (Echo-high), meaning it is inside the region with the highest average opinion. Here, the engagement range shrinks to 0.25–0.75, while its maximum value increases up to 100. Finally, the last plot represents engagement over a uniformly distributed network with high modularity and low homophily, where the source node is in the low-opinion echo chamber (Echo-low). In this case, engagement peaks around the dominant opinion within the echo chamber.
Appendix B Simulated Against Real-World Engagement
@@
We validate the reliability of our propagation model by comparing the estimated engagement over the Brexit real network with the actual number of retweets generated by the most engaging content in the dataset. Figure 14 presents the corresponding regression line, computed using the Random Sample Consensus (RANSAC) method Cantzler ([n. d.]). Although our propagation protocol tends to overestimate engagement, we observe a positive correlation between the number of active users predicted by our model ($|\mathcal{A}|$) and the actual retweet count. This correlation supports the validity of our findings on the real network.